Write the equation of a line parallel to the given line but passing through the given point.
y= -1/3x -4
pass through points are -6 and -3

Answer by Cromlix(4381)

Lines that are parallel share the same
gradient/slope/m
y = -1/3x - 4
has a gradient of -1/3
Using y - b = m(x - a)
Where a = -6, b = -3 and m = -1/3
y -(-3) = -1/3)x -(-6)
y + 3 = -1/3x - 2
y = -1/3x - 2 - 3
y = -1/3x - 5
Hope this helps.
:-)